OpenStudy (anonymous):

What would be the steps for to solve 2x + 5 = 31 for x.

OpenStudy (anonymous):

2x + 5 − 5 = 31 − 5 2x + 0 = 26 2x = 26 2x ÷ 2 = 26 ÷ 2 x = 13.

OpenStudy (anonymous):

First, subtract 5 from both sides so the five becomes 0 and the 31 becomes 26. They you are left with 2x=26. Then divide by 2 on both sides so x is just left on the left and 13 on the other. x=13

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

Any time

OpenStudy (anonymous):

i gave you the right asnwer and you never rated me.lmaoo what a trip yeoo